# Quantum State Preparation with Optimal T-Count

### v1

@@ 4.1 State preparation
An operator $P$ is an $n$-qubit Pauli iff $P=i^{b}P^{\prime}$ for some $b\in\left\{0,1,2,3\right\}$ and $P^{\prime}\in\{I,X,Y,Z\}^{\otimes n}$. In addition, if $b\in\left\{0,2\right\}$, then $P$ is a Hermitian Pauli. An $n$-qubit Pauli postselection is specified by an $n$-qubit Hermitian Pauli operator $P$. On input state $|\phi\rangle$, the postselection produces a normalized state $|\psi\rangle$ of $|\phi\rangle$ projected to the $+1$ eigenspace of $P$, i.e., $|\psi\rangle$ is proportional to $(I+P)|\phi\rangle$, denoted by $|\psi\rangle\propto(I+P)|\phi\rangle$. More precisely, $|\psi\rangle=\frac{(I+P)|\phi\rangle}{\left\|(I+P)|\phi\rangle\right\|}$ and it is defined arbitrarily if $(I+P)|\phi\rangle$ becomes zero. Note that Pauli postselection is not a linear map.
@@
A Clifford circuit with Pauli postselection and copies of the single-qubit magic state $|T\rangle=\frac{1}{\sqrt{2}}(|0\rangle+e^{i\pi/4}|1\rangle)$ can perform ${T}$ gates by gate injection. [^16] In particular, we can implement a ${T}$ gate using the fact that $(I\otimes\langle 0|)\mathrm{CNOT}|\psi\rangle|T\rangle=\frac{1}{\sqrt{2}}T|\psi\rangle$ for any single qubit $|\psi\rangle$. In the following we shall consider state preparation circuits of the following form: $\mathcal{C}$ is a Clifford circuit with $m$ Pauli postselections, $n$ input qubits, and $a$ ancillas such that
@@
$|\psi_{\text{out}}\rangle|0^{t}\rangle|0^{a}\rangle\propto C_{m+1}M_{m}C_{m}\cdots M_{1}C_{1}\left(|\phi_{\text{in}}\rangle|{T}\rangle^{\otimes t}|0^{a}\rangle\right),$ (52)
@@
The following Claim shows that, if there is an adaptive Clifford+${T}$ circuit $\mathcal{A}$ that approximately prepares a given state with expected ${T}$-count $\mathcal{T}^{0}(\mathcal{A})$, then there is a Clifford circuit with Pauli postselections and $2\mathcal{T}^{0}(\mathcal{A})$ magic states that also approximately prepares the state. This allows us to shift our focus to the model of Clifford circuits with Pauli postselection. The proof uses an averaging argument similar to the one used in [5, Lemma 5.4].
@@
Assume $\mathcal{A}$ is an adaptive Clifford+${T}$ circuit that prepares an $n$-qubit state $|\psi\rangle$ up to error $\varepsilon$ in trace distance. Let $t=2\cdot\mathcal{T}^{0}(\mathcal{A})$. Then for some integers $m,a\geq 0$, there exists a Clifford circuit $\mathcal{C}$ with $m$ Pauli postselections and $a$ ancillas such that $\mathcal{C}\left(|0^{n}\rangle|{T}\rangle^{\otimes t}|0^{a}\rangle\right)=|\phi\rangle|0^{t+a}\rangle$ and $\frac{1}{2}\left\|\left|\psi\right\rangle\left\langle\psi\right|-\left|\phi\right\rangle\left\langle\phi\right|\right\|_{1}\leq\sqrt{6\varepsilon}$.
@@
where $p_{r}(\rho)$ is a distribution. In addition, $K_{r}$ uses $\mathcal{T}_{r}$ ${T}$ gates and
$\sigma_{r}(\rho)=\left|\pi_{r}\right\rangle\left\langle\pi_{r}\right|,\quad\text{where}\quad|\pi_{r}\rangle\propto K_{r}\left(|0^{n}\rangle|0^{a}\rangle\right)\quad\text{is a pure state.}$. (54)
We will find some $r$ such that $|\pi_{r}\rangle$ is close to $|\psi\rangle|0^{a}\rangle$ and $K_{r}$ uses a small number of ${T}$ gates.
@@
where we use the inequality $1-\sqrt{F}\leq\varepsilon$ (see e.g., [50, Theorem 9.3.1]). Plugging this into Equation 53 and Equation 54, we have
@@
we know that the fidelity between $|\phi\rangle$ and $|\psi\rangle$ is at least $1-6\varepsilon$. This implies that the trace distance between $|\phi\rangle$ and $|\psi\rangle$ is at most $\sqrt{6\varepsilon}$ (see e.g., [50, Theorem 9.3.1]).
Finally we observe that $|\phi\rangle|0^{a}\rangle$ is obtained from $|\pi_{r}\rangle$ followed by postselecting the last $a$ qubits being all-zero. This means we can construct $|\phi\rangle|0^{a}\rangle$ by applying $K_{r}$ and the above postselection on $|0^{n}\rangle|0^{a}\rangle$. This process uses $t=2\cdot\mathcal{T}^{0}(\mathcal{A})$ ${T}$ gates, [^17] Though $\mathcal{T}_{r}\leq 2\cdot\mathcal{T}(\mathcal{A})$, we can always ensure that there are exactly $t$ ${T}$ gates (say by applying dummy ${T}$ gates acting on one of the ancillas in the $|0\rangle$ state at the beginning of the circuit; note $T|0\rangle=|0\rangle$). each of which can be implemented using a magic state $|T\rangle$ and gate injection. This gives a Clifford circuit $\mathcal{C}$ with Pauli postselection and input state $|0^{n}\rangle|{T}\rangle^{\otimes t}|0^{a}\rangle$ which prepares $|\phi\rangle|0^{a}\rangle$. ∎
Given Claim 4.5, we now focus on Clifford+${T}$ circuits with Pauli postselections and ancillas. To complete the proof of Theorem 4.1, it is sufficient to prove the following proposition.
@@
The first lower bound of $\Omega(\log(1/\varepsilon))$ holds even for single-qubit state preparation. This is proved by Beverland, Campbell, Howard, and Kliuchnikov [5, Lemma 5.9]. The second ${T}$-count lower bound is a counting argument as in [31]: the number of distinct circuits should be at least the number of distinguishable states in question. The latter quantity follows directly from standard sphere packing bounds for $\ell_{2}$ balls (see e.g., [48]).
@@
Assume $\varepsilon<1/8$. This ensures that the trace distance between $|\phi\rangle$ and $|\psi\rangle$ is at most $1/8$ (see e.g., [50, Section 9]). Then the $\Omega(\log(1/\varepsilon))$ lower bound follows from [5, Lemma 5.9].
@@ 4.2 Unitary synthesis
Now we turn to the unitary synthesis ${T}$-count lower bounds (Theorem 4.2 and Theorem 4.3). We will reduce them to state preparation tasks, by considering the action of an adaptive Clifford+${T}$ circuit on the maximally entangled state and comparing this with the Choi state corresponding to the unitary $U$ of interest.
@@
Assume an adaptive Clifford+${T}$ circuit $\mathcal{A}$, using $a$ ancillas, implements an $n$-qubit unitary $U$ up to error $\varepsilon$ in diamond distance. Then $(I\otimes\mathcal{A})(\left|\iota\right\rangle\left\langle\iota\right|)$ is $\varepsilon$-close in trace distance to $\left|\iota_{U}\right\rangle\left\langle\iota_{U}\right|\otimes\left|0^{a}\right\rangle\left\langle 0^{a}\right|$.
@@
Assume an adaptive Clifford+${T}$ circuit $\mathcal{A}$ implements an $n$-qubit unitary $U$ up to error $\varepsilon$ in the diamond distance. Let $t=2\cdot\mathcal{T}(\mathcal{A})$. Then for some integers $m,a\geq 0$, there exists a Clifford circuit $\mathcal{C}$ with $m$ Pauli postselections and $a$ ancillas such that $\mathcal{C}\left(|0^{2n}\rangle|{T}\rangle^{\otimes t}|0^{a}\rangle\right)=|\phi\rangle|0^{t+a}\rangle$ and $\frac{1}{2}\left\|\left|\iota_{U}\right\rangle\left\langle\iota_{U}\right|-\left|\phi\right\rangle\left\langle\phi\right|\right\|_{1}\leq\sqrt{6\varepsilon}$.
@@
By Fact 4.10, we know that $(I\otimes\mathcal{A})(\left|\iota\right\rangle\left\langle\iota\right|)$ is $\varepsilon$-close in trace distance to $\left|\iota_{U}\right\rangle\left\langle\iota_{U}\right|\otimes\left|0^{a}\right\rangle\left\langle 0^{a}\right|$ for some integer $a\geq 0$. Then following the proof of Claim 4.5 (but replacing the all-zeros input state with $\rho=\left|\iota\right\rangle\left\langle\iota\right|$ and the expected ${T}$-count $\mathcal{T}^{0}(\mathcal{A})$ by the worst-case expected ${T}$-count $\mathcal{T}(\mathcal{A})$), we obtain a Clifford circuit $\mathcal{C}^{\prime}$ with Pauli postselection such that
@@
In light of Fact 4.14, we consider the diagonal unitaries $D=D_{j}$ for each $j\in[N]$ from Fact 4.15. Then by Lemma 4.8, we obtain a canonical form for the $2n+t$-qubit Clifford circuit with Pauli postselection that approximately prepares $|\iota_{D_{j}}\rangle$ for each $j$. But there are at most $2^{O(n^{2}+t^{2})}$ different canonical forms on $2n+t$ qubits. By Fact 4.14 and Fact 4.15, each $|\iota_{D_{j}}\rangle$ gives a different circuit. Thus $2^{O(n^{2}+t^{2})}\geq(1/\varepsilon)^{\Theta(2^{n})}$, which gives $t=\Omega\left(\sqrt{2^{n}\log(1/\varepsilon)}\right)$ since $\varepsilon<1/8$. ∎
@@ Acknowledgments
We thank Ryan Babbush, Dominic Berry, Michael Beverland, Oscar Higgott, Tanuj Khattar, Guang Hao Low, Dmitri Maslov, Luke Shaeffer, and Rolando Somma for helpful discussions. We thank Tanuj Khattar for pointing out that the analysis of the algorithm in [31] can be improved. KW wants to thank Daniel Grier, Jiaqing Jiang, Saeed Mehraban, and Anurudh Peduri for relevant references. DG is a CIFAR fellow in the quantum information science program. KW is supported by a Sloan Research Fellowship and NSF CAREER Award CCF-2145474, and this work was done while at Google.
@@ Appendix A Canonical form of Clifford circuits with magic states and Pauli postselection
For a state $|\tau\rangle$, its stabilizer group $\mathsf{Stab}(|\tau\rangle)$ is the set of Pauli operators $P$ satisfying $P|\tau\rangle=|\tau\rangle$. For a set $\Lambda$ of states, we define its stabilizer group as $\mathsf{Stab}(\Lambda)=\bigcap_{|\tau\rangle\in\Lambda}\mathsf{Stab}(|\tau\rangle)$. Note that the stabilizer group is a set of commutative Hermitian Pauli operators and does not contain $-I$. We also remark that the size of a stabilizer group is an integer power of two.
@@
Intuitively, Theorem A.2 says that the number of ancillary qubits and the number of Pauli postselections are upper bounded by the number of ${T}$ gates.
@@
Let $P,Q$ be two anticommuting Hermitian Pauli operators. Then $-iPQ$ is Hermitian Pauli and $\frac{I+PQ}{\sqrt{2}}$ is Clifford.
@@
By Item 2 of Lemma A.4, we know that each $Q_{j}$ commutes with $\mathsf{Stab}\left(\left\{|\phi_{\lambda}\rangle|{T}\rangle^{\otimes t}|0^{a}\rangle\right\}_{\lambda\in\mathcal{S}}\right)$, which contains all ${Z}$ rotations on the last $a$-qubits. Since $Q_{j}$ is a Hermitian Pauli, we have
$Q_{j}=P_{j}\otimes{Z}(v_{j})\quad\text{for some $(n+t)$-qubit Hermitian Pauli $P_{j}$ and some $v_{j}\in\{0,1\}^{a}$.}$ (85)
Here, ${Z}(v_{j})$ applies the single-qubit ${Z}$ gate on the coordinates selected by $v_{j}$. Define $M_{j}=I+P_{j}$. Then for any $(n+t)$-qubit state $|\rho\rangle$, we have
@@ A.1 Eliminating ancillas in Clifford circuits
Let $r=\log\left|\mathsf{Stab}(\left\{|\phi_{\lambda}\rangle\right\})\right|$. Since $\mathsf{Stab}(\left\{|\phi_{\lambda}\rangle\right\})$ is a commutative subgroup of the Pauli group and does not contain $-I$, there exists a Clifford $C_{\phi}$ conjugating $\mathsf{Stab}(\left\{|\phi_{\lambda}\rangle\right\})$ to the group of ${Z}$ operators on the last $r$ qubits [14], which means
$|\phi_{\lambda}\rangle=C_{\phi}\left(|\phi_{\lambda}^{\prime}\rangle|0^{r}\rangle\right)\quad\text{holds for every $\lambda\in\mathcal{S}$ and some $(n-r)$-qubit state $|\phi_{\lambda}^{\prime}\rangle$.}$ (88)
@@
$|\psi_{\lambda}\rangle=C_{\psi}\left(|\psi_{\lambda}^{\prime}\rangle|0^{r}\rangle\right)\quad\text{holds for every $\lambda\in\mathcal{S}$ and some $(n-r)$-qubit state $|\psi_{\lambda}^{\prime}\rangle$.}$ (89)
@@
We will make another simplification to make sure that $C^{\prime\prime}$ commutes with the single-qubit ${Z}$ gate on the last qubit, which will help factorize $C^{\prime\prime}$ and apply induction. To this end, we observe that $\mathsf{Stab}(\left\{|\phi_{\lambda}^{\prime}\rangle|0^{r+m}\rangle\right\})=\mathsf{Stab}(\left\{|\psi_{\lambda}^{\prime}\rangle|0^{r+m}\rangle\right\})$ equals the Pauli ${Z}$ group on the last $r+m$ qubits. Hence by Equation 90, for the single-qubit ${Z}$ gate on the last qubit we know that
@@
Let $D$ be an $(r+m)$-qubit CNOT circuit reversing this mapping, i.e., $D{Z}(s)D^{\dagger}$ equals the single-qubit ${Z}$ gate on the last qubit. Define Clifford $C^{\prime\prime\prime}=(I_{n-r}\otimes D)C^{\prime\prime}$. Then we have the commutativity condition:
@@
Assume $C$ is a $K$-qubit Clifford unitary commuting with the single-qubit ${Z}$ gate on the last qubit. Then $C=C_{0}\otimes\left|0\right\rangle\left\langle 0\right|+C_{1}\otimes\left|1\right\rangle\left\langle 1\right|$, where $C_{0},C_{1}$ are $(K-1)$-qubit Clifford unitaries.
@@
where $C_{0}$ is an $(n+m-1)$-qubit Clifford. Furthermore, by Equation 93, we have
@@
Iterating the above process to ensure $C_{0}$ commutes with the single-qubit ${Z}$ gate on the last qubit, we can repeatedly trim off the trailing ancillas until we obtain

### v2

@@ 4.1 State preparation
An operator $P$ is an $n$-qubit Pauli iff $P=i^{b}P^{\prime}$ for some $b\in\left\{0,1,2,3\right\}$ and $P^{\prime}\in\{I,X,Y,Z\}^{\otimes n}$. In addition, if $b\in\left\{0,2\right\}$, then $P$ is a Hermitian Pauli. An $n$-qubit Pauli postselection is specified by an $n$-qubit Hermitian Pauli operator $P$. On input state $|\phi\rangle$, the postselection produces a normalized state $|\psi\rangle$ of $|\phi\rangle$ projected to the +1+1 eigenspace of $P$, i.e., $|\psi\rangle$ is proportional to $(I+P)|\phi\rangle$, denoted by $|\psi\rangle\propto(I+P)|\phi\rangle$. More precisely, $|\psi\rangle=\frac{(I+P)|\phi\rangle}{\left\|(I+P)|\phi\rangle\right\|}$ and it is defined arbitrarily if $(I+P)|\phi\rangle$ becomes zero. Note that Pauli postselection is not a linear map.
@@
A Clifford circuit with Pauli postselection and copies of the single-qubit magic state $|T\rangle=\frac{1}{\sqrt{2}}(|0\rangle+e^{i\pi/4}|1\rangle)$ can perform T{T} gates by gate injection. [^24] In particular, we can implement a T{T} gate using the fact that $(I\otimes\langle 0|)\mathrm{CNOT}|\psi\rangle|T\rangle=\frac{1}{\sqrt{2}}T|\psi\rangle$ for any single qubit $|\psi\rangle$. In the following we shall consider state preparation circuits of the following form: $\mathcal{C}$ is a Clifford circuit with $m$ Pauli postselections, $n$ input qubits, and $a$ ancillas such that
@@
$|\phi_{\text{out}}\rangle|0^{t}\rangle|0^{a}\rangle\propto C_{m+1}M_{m}C_{m}\cdots M_{1}C_{1}\left(|\phi_{\text{in}}\rangle|{T}\rangle^{\otimes t}|0^{a}\rangle\right),$ (52)
@@
The following Claim shows that, if there is an adaptive Clifford+T{T} circuit $\mathcal{A}$ that approximately prepares a given state with expected T{T}-count $\mathcal{T}^{0}(\mathcal{A})$, then there is a Clifford circuit with Pauli postselections and $2\mathcal{T}^{0}(\mathcal{A})$ magic states that also approximately prepares the state. This allows us to shift our focus to the model of Clifford circuits with Pauli postselection. The proof uses an averaging argument similar to the one used in [5, Lemma 5.4].
@@
Assume $\mathcal{A}$ is an adaptive Clifford+T{T} circuit that prepares an $n$-qubit state $|\psi\rangle$ up to error $\varepsilon$ in trace distance. Let $t=2\cdot\mathcal{T}^{0}(\mathcal{A})$. Then for some integers $m,a\geq 0$, there exists a Clifford circuit $\mathcal{C}$ with $m$ Pauli postselections and $a$ ancillas such that $\mathcal{C}\left(|0^{n}\rangle|{T}\rangle^{\otimes t}|0^{a}\rangle\right)=|\phi\rangle|0^{t+a}\rangle$ and $\frac{1}{2}\left\|\left|\psi\right\rangle\left\langle\psi\right|-\left|\phi\right\rangle\left\langle\phi\right|\right\|_{1}\leq\sqrt{6\varepsilon}$.
@@
where $p_{r}(\rho)$ is a distribution. In addition, $K_{r}$ uses $\mathcal{T}_{r}$ T{T} gates and
$\sigma_{r}(\rho)=\left|\pi_{r}\right\rangle\left\langle\pi_{r}\right|,\quad\text{where}\quad|\pi_{r}\rangle\propto K_{r}\left(|0^{n}\rangle|0^{a}\rangle\right)\quad\text{is a pure state.}$ (54)
We will find some $r$ such that $|\pi_{r}\rangle$ is close to $|\psi\rangle|0^{a}\rangle$ and $K_{r}$ uses a small number of T{T} gates.
@@
where we use the inequality $1-\sqrt{F}\leq\varepsilon$ (see e.g., [51, Theorem 9.3.1]). Plugging this into Equation 53 and Equation 54, we have
@@
we know that the fidelity between $|\phi\rangle$ and $|\psi\rangle$ is at least $1-6\varepsilon$. This implies that the trace distance between $|\phi\rangle$ and $|\psi\rangle$ is at most $\sqrt{6\varepsilon}$ (see e.g., [51, Theorem 9.3.1]).
Finally we observe that $|\phi\rangle|0^{a}\rangle$ is obtained from $|\pi_{r}\rangle$ followed by postselecting the last $a$ qubits being all-zero. This means we can construct $|\phi\rangle|0^{a}\rangle$ by applying $K_{r}$ and the above postselection on $|0^{n}\rangle|0^{a}\rangle$. This process uses $t=2\cdot\mathcal{T}^{0}(\mathcal{A})$ T{T} gates, [^25] Though $\mathcal{T}_{r}\leq 2\cdot\mathcal{T}(\mathcal{A})$, we can always ensure that there are exactly $t$ T{T} gates (say by applying dummy T{T} gates acting on one of the ancillas in the $|0\rangle$ state at the beginning of the circuit; note $T|0\rangle=|0\rangle$). each of which can be implemented using a magic state $|T\rangle$ and gate injection. This gives a Clifford circuit $\mathcal{C}$ with Pauli postselection and input state $|0^{n}\rangle|{T}\rangle^{\otimes t}|0^{a}\rangle$ which prepares $|\phi\rangle|0^{a}\rangle$. ∎
Given Claim 4.5, we now focus on Clifford+T{T} circuits with Pauli postselections and ancillas. To complete the proof of Theorem 4.1, it is sufficient to prove the following proposition.
@@
The first lower bound of $\Omega(\log(1/\varepsilon))$ holds even for single-qubit state preparation. This is proved by Beverland, Campbell, Howard, and Kliuchnikov [5, Lemma 5.9]. The second T{T}-count lower bound is a counting argument as in [31]: the number of distinct circuits should be at least the number of distinguishable states in question. The latter quantity follows directly from standard sphere packing bounds for $\ell_{2}$ balls (see e.g., [49]).
@@
Assume $\varepsilon<1/8$. This ensures that the trace distance between $|\phi\rangle$ and $|\psi\rangle$ is at most $1/8$ (see e.g., [51, Section 9]). Then the $\Omega(\log(1/\varepsilon))$ lower bound follows from [5, Lemma 5.9].
@@ 4.2 Unitary synthesis
Now we turn to the unitary synthesis T{T}-count lower bounds (Theorem 4.2 and Theorem 4.3). We will reduce them to state preparation tasks, by considering the action of an adaptive Clifford+T{T} circuit on the maximally entangled state and comparing this with the Choi state corresponding to the unitary $U$ of interest.
@@
Assume an adaptive Clifford+T{T} circuit $\mathcal{A}$, using $a$ ancillas, implements an $n$-qubit unitary $U$ up to error $\varepsilon$ in diamond distance. Then $(I\otimes\mathcal{A})(\left|\iota\right\rangle\left\langle\iota\right|)$ is $\varepsilon$-close in trace distance to $\left|\iota_{U}\right\rangle\left\langle\iota_{U}\right|\otimes\left|0^{a}\right\rangle\left\langle 0^{a}\right|$.
@@
Assume an adaptive Clifford+T{T} circuit $\mathcal{A}$ implements an $n$-qubit unitary $U$ up to error $\varepsilon$ in the diamond distance. Let $t=2\cdot\mathcal{T}(\mathcal{A})$. Then for some integers $m,a\geq 0$, there exists a Clifford circuit $\mathcal{C}$ with $m$ Pauli postselections and $a$ ancillas such that $\mathcal{C}\left(|0^{2n}\rangle|{T}\rangle^{\otimes t}|0^{a}\rangle\right)=|\phi\rangle|0^{t+a}\rangle$ and $\frac{1}{2}\left\|\left|\iota_{U}\right\rangle\left\langle\iota_{U}\right|-\left|\phi\right\rangle\left\langle\phi\right|\right\|_{1}\leq\sqrt{6\varepsilon}$.
@@
By Fact 4.10, we know that $(I\otimes\mathcal{A})(\left|\iota\right\rangle\left\langle\iota\right|)$ is $\varepsilon$-close in trace distance to $\left|\iota_{U}\right\rangle\left\langle\iota_{U}\right|\otimes\left|0^{a}\right\rangle\left\langle 0^{a}\right|$ for some integer $a\geq 0$. Then following the proof of Claim 4.5 (but replacing the all-zeros input state with $\rho=\left|\iota\right\rangle\left\langle\iota\right|$ and the expected T{T}-count $\mathcal{T}^{0}(\mathcal{A})$ by the worst-case expected T{T}-count $\mathcal{T}(\mathcal{A})$), we obtain a Clifford circuit $\mathcal{C}^{\prime}$ with Pauli postselection such that
@@
In light of Fact 4.14, we consider the diagonal unitaries $D=D_{j}$ for each $j\in[N]$ from Fact 4.15. Then by Lemma 4.8, we obtain a canonical form for the $(2n+t)$-qubit Clifford circuit with Pauli postselection that approximately prepares $|\iota_{D_{j}}\rangle$ for each $j$. But there are at most $2^{O(n^{2}+t^{2})}$ different canonical forms on $2n+t$ qubits. By Fact 4.14 and Fact 4.15, each $|\iota_{D_{j}}\rangle$ gives a different circuit. Thus $2^{O(n^{2}+t^{2})}\geq(1/\varepsilon)^{\Theta(2^{n})}$, which gives $t=\Omega\left(\sqrt{2^{n}\log(1/\varepsilon)}\right)$ since $\varepsilon<1/8$. ∎
@@ Acknowledgments
We thank Ryan Babbush, Dominic Berry, Michael Beverland, Oscar Higgott, Tanuj Khattar, Guang Hao Low, Dmitri Maslov, Luke Shaeffer, and Rolando Somma for helpful discussions. We thank Tanuj Khattar for pointing out that the analysis of the algorithm in [31] can be improved. We thank anonymous reviewers for helpful comments. KW wants to thank Daniel Grier, Jiaqing Jiang, Saeed Mehraban, and Anurudh Peduri for relevant references.
DG is a CIFAR fellow in the quantum information science program. KW is supported by the National Science Foundation under Grant No. DMS-2424441, and by the IAS School of Mathematics; this work was done while at Google.
@@ Appendix A Canonical form of Clifford circuits with magic states and Pauli postselection
For a state $|\tau\rangle$, its stabilizer group $\mathsf{Stab}(|\tau\rangle)$ is the set of Pauli operators $P$ satisfying $P|\tau\rangle=|\tau\rangle$. For a set $\Lambda$ of states, we define its stabilizer group as $\mathsf{Stab}(\Lambda)=\bigcap_{|\tau\rangle\in\Lambda}\mathsf{Stab}(|\tau\rangle)$. Note that the stabilizer group is a set of commutative Hermitian Pauli operators and does not contain I-I. We also remark that the size of a stabilizer group is an integer power of two.
@@
Intuitively, Theorem A.2 says that the number of ancillary qubits and the number of Pauli postselections are upper bounded by the number of T{T} gates.
@@
Let $P,Q$ be two anticommuting Hermitian Pauli operators. Then iPQ-iPQ is Hermitian Pauli and $\frac{I+PQ}{\sqrt{2}}$ is Clifford.
@@
By Item 2 of Lemma A.4, we know that each $Q_{j}$ commutes with $\mathsf{Stab}\left(\left\{|\phi_{\lambda}\rangle|{T}\rangle^{\otimes t}|0^{a}\rangle\right\}_{\lambda\in\mathcal{S}}\right)$, which contains all Z{Z} rotations on the last $a$-qubits. Since $Q_{j}$ is a Hermitian Pauli, we have
[…]
Here, ${Z}(v_{j})$ applies the single-qubit Z{Z} gate on the coordinates selected by $v_{j}$. Define $M_{j}=I+P_{j}$. Then for any $(n+t)$-qubit state $|\rho\rangle$, we have
@@ A.1 Eliminating ancillas in Clifford circuits
Let $r=\log\left|\mathsf{Stab}(\left\{|\phi_{\lambda}\rangle\right\})\right|$. Since $\mathsf{Stab}(\left\{|\phi_{\lambda}\rangle\right\})$ is a commutative subgroup of the Pauli group and does not contain I-I, there exists a Clifford $C_{\phi}$ conjugating $\mathsf{Stab}(\left\{|\phi_{\lambda}\rangle\right\})$ to the group of Z{Z} operators on the last $r$ qubits [14], which means
[…]
@@
[…]
@@
We will make another simplification to make sure that $C^{\prime\prime}$ commutes with the single-qubit Z{Z} gate on the last qubit, which will help factorize $C^{\prime\prime}$ and apply induction. To this end, we observe that $\mathsf{Stab}(\left\{|\phi_{\lambda}^{\prime}\rangle|0^{r+m}\rangle\right\})=\mathsf{Stab}(\left\{|\psi_{\lambda}^{\prime}\rangle|0^{r+m}\rangle\right\})$ equals the Pauli Z{Z} group on the last $r+m$ qubits. Hence by Equation 90, for the single-qubit Z{Z} gate on the last qubit we know that
@@
Let $D$ be an $(r+m)$-qubit CNOT circuit reversing this mapping, i.e., $D{Z}(s)D^{\dagger}$ equals the single-qubit Z{Z} gate on the last qubit. Define Clifford $C^{\prime\prime\prime}=(I_{n-r}\otimes D)C^{\prime\prime}$. Then we have the commutativity condition:
@@
Assume $C$ is a $K$-qubit Clifford unitary commuting with the single-qubit Z{Z} gate on the last qubit. Then $C=C_{0}\otimes\left|0\right\rangle\left\langle 0\right|+C_{1}\otimes\left|1\right\rangle\left\langle 1\right|$, where $C_{0},C_{1}$ are (K1)(K-1)-qubit Clifford unitaries.
@@
where $C_{0}$ is an (n+m1)(n+m-1)-qubit Clifford. Furthermore, by Equation 93, we have
@@
Iterating the above process to ensure $C_{0}$ commutes with the single-qubit Z{Z} gate on the last qubit, we can repeatedly trim off the trailing ancillas until we obtain